OpenStudy (anonymous):

what measure of angle A D=79 BC=73 i kNow B=101 but i am confused about A sense I am not sure what angle C is D+b=180 A+C=180

OpenStudy (anonymous):

 abdc

jhonyy9 (jhonyy9):

so you know that A+B+C+D=360 try getting from what you know begined with this

OpenStudy (kaederfds):

What is the meaning of "BC=73" ?

OpenStudy (anonymous):

probably the arc

jhonyy9 (jhonyy9):

D=79 B=101 --- A+B=180 --- A+101=180 --- A=180-101 --- A=79 - from this angle measures result that this draw not is right sure

jhonyy9 (jhonyy9):

change sign B with D for being right measure of angles

OpenStudy (anonymous):

thanks I know why its not correct

jhonyy9 (jhonyy9):

so now you can solve it right sure ?

OpenStudy (anonymous):

yes its actually asking for arc AB i got it thank you

jhonyy9 (jhonyy9):

yw good luck bye